# An Image is Worth 1/2 Tokens After Layer 2: Plug-and-Play Inference Acceleration for Large Vision-Language Models

### 11 Mar 2024 v1

@@ 1 Introduction
To address the problem, we assume a plausible explanation is that the high redundancy in visual signals leads to the aggregation of image-related, instruction-specific features onto certain “anchor” tokens through the self-attention mechanism in the shallow layers. Notably, these anchor tokens are rarely image tokens. In deep layers, attentions are focused on those anchor tokens, leading to significantly reduced attention on the image tokens themselves.
@@ 3.3 Results
Figure 4: The attention maps during the decoding process of one model response for LLaVA1.5-7B. We can see that in the bottom layer, attention distributes relatively smooth across different type of tokens. In the the deep layers, above from local attention, the attention scores are aggregated to system prompt, instruction and output tokens and attention over image tokens is rather sparse.
@@ 3.4 Insights
From the attention visualization results, we can see that in shallow layer, the attention scores distribute more smoothly across different tokens. While in deep layer, there are vertical strong lines (in the system prompt) that takes up most of attention scores. The existence of vertical strong line shows that there are some input tokens that consistently received high attention during the whole decoding process. This also explains the highly imbalanced attention efficiencies in our statistics: A small portion of anchor tokens aggregate the information from all input tokens and the model much favors to attend to those anchor tokens in deep layers. Much information of image tokens are aggregated to non-image tokens such as system prompt, which leads to the severe inefficient visual attention in LVLMs. Our findings also align with the information flow of Large Language Model found in  Wang et al. (2023).
@@ 5.3 Main Results
(d) w/ FastV (system prompt) 89.2 64.3 69.2 33.8
(e) w/ FastV (prune first half system prompt) 17.5 27.8 Failed Failed
(f) w/ FastV (instruction) 77.3 50.1 56.5 29.5
@@ 5.4 Ablation Studies
As we find that LVLM and LLM bear similar attention patterns that the head tokens play an important role during infernce, we wonder that whether the same attention optimization technique curated for LLM can be transfered to LVLM. Following the methodology outlined in StreamingLLM Xiao et al. (2023), we implemented an attention optimization scheme that incorporates an attention sink for the first four tokens and local attention for the preceding four tokens throughout the decoding process. The experiment’s outcome, presented in line (g) of Table 5, indicates a substantial degradation in LVLM’s performance when this specific attention pattern is applied. This suggests a fundamental difference in how image tokens, as opposed to text tokens, contribute to the information processing within LLMs. This observation highlights the necessity for further investigation into the unique behaviors of image tokens within the LVLM framework, suggesting that the direct application of LLM attention/KV-Cache optimization mechanisms to LVLMs requires adaptation to accommodate the distinct characteristics of visual information.

### 25 Mar 2024 v2

@@ 1 Introduction
To address the problem, we assume a plausible explanation is that the high redundancy in visual signals leads to the aggregation of image-related, instruction-specific features onto certain “anchor” tokens through the self-attention mechanism in the shallow layers. Notably, these anchor tokens might not be image tokens. In deep layers, attentions are focused on those anchor tokens, leading to significantly reduced attention on the image tokens themselves.
@@ 3.3 Results
Figure 4: The attention maps during the decoding process of one model response for LLaVA1.5-7B. We can see that in the bottom layer, attention distributes relatively smooth across different type of tokens. In the the deep layers, above from local attention, the attention scores are aggregated to system prompt, instruction and output tokens. The attention over image tokens is rather sparse.
@@ 3.4 Insights
From the attention visualization results, we can see that in shallow layer, the attention scores distribute more smoothly across different tokens. While in deep layer, there are vertical strong lines (in the system prompt) that takes up most of attention scores. The existence of vertical strong line shows that there are some input tokens that consistently received high attention during the whole decoding process. This also explains the highly imbalanced attention efficiencies in our statistics: A small portion of anchor tokens draw the most attention and the model much favors to attend to those anchor tokens in deep layers. Much attention is aggregated to the beginning system prompt, which leads to the severe inefficient visual attention in LVLMs. Our findings also align with the attention sink phenomena of Large Language Model found in  Xiao et al. (2023).
@@ 5.3 Main Results
(d) w/ FastV (K=2, R=50%, on system prompt) 89.2 64.3 69.2 33.8
(e) w/ FastV (K=2, Prune first half system prompt) 17.5 27.8 Failed Failed
(f) w/ FastV (K=2, R=50%, on instruction) 77.3 50.1 56.5 29.5
@@ 5.4 Ablation Studies
As we find that LVLM and LLM bear similar attention patterns that the head tokens play an important role during inference, we wonder that whether the same attention optimization technique curated for LLM can be transfered to LVLM. Following the methodology outlined in StreamingLLM Xiao et al. (2023), we implemented an attention optimization scheme that incorporates an attention sink for the first four tokens and keep all local attention (context size=4) for the preceding four tokens throughout the decoding process. The experiment’s outcome, presented in line (g) of Table 5, indicates a substantial degradation in LVLM’s performance when this specific attention pattern is applied. This suggests a fundamental difference in how image tokens, as opposed to text tokens, contribute to the information processing within LLMs. This observation highlights the necessity for further investigation into the unique behaviors of image tokens within the LVLM framework, suggesting that the direct application of LLM attention/KV-Cache optimization mechanisms to LVLMs requires adaptation to accommodate the distinct characteristics of visual information.